# An FPGA-based Versatile and Digitalized Method for Pulse Laser Repetition Frequency Locking

### v1

@@ Abstract
This paper introduces a novel method for locking the repetition frequency of pulse lasers, adaptable to different frequencies, offering significant improvements in system integration and measurement accuracy. The method consists of two primary components: an error amplification module (EAM) and a digital frequency locking module (DFLM) based on FPGA. The EAM integrates a configurable frequency generator (CFG), a configurable frequency multiplier (CFM) and a mixer, to process the laser pulse alongside a high-stability reference source, such as an atomic clock. By employing frequency multiplication and mixing, the EAM amplifies the laser’s frequency error and performs frequency down-conversion, enhancing measurement sensitivity and reducing the hardware requirements of the back-end. The CFG, implemented on a phase-locked loop (PLL) chip, allows for parameter adjustments to accommodate various laser frequencies. The DFLM processes the output from the EAM using a high-speed, ADC-based dual-mixer time-difference (DMTD) method to precisely measure frequency errors. A digital proportional-integral-derivative (PID) controller then provides feedback to achieve accurate frequency locking. To evaluate the proposed method, an FPGA-based electronic system was developed and tested. In laboratory experiment with a custom-built femtosecond fiber laser, the system demonstrated robust locking of the repetition rate, achieving an Allan deviation improvement from $1.51\times 10^{-7}$ to $1.12\times 10^{-12}$ at a gate time of 10 s. Further testing with a voltage-controlled oscillator (VCO) confirmed a long-term stability of $9.58\times 10^{-14}@10s$.
@@
Field programmable gate array (FPGA), laser frequency locking, frequency error amplifier, digital electronic circuits, dual-mixer time-difference (DMTD).
@@ I Introduction
Pulse lasers characterized by narrow-lines and high frequency stability are essential components in optical frequency combs. They have widespread applications in both industrial settings and scientific research, including high-precision measurement[1][2], optical communication[3][4], high-resolution optical spectroscopy[5][6], etc. Optical frequency combs are advanced tools that generate millions of mutually coherent and evenly spaced modes, enabling precise phase-coherent connections across the optical, microwave, and terahertz frequency ranges. These frequency combs are essential for a variety of applications and fundamental scientific research[7][8]. However, optic combs have been limited by various factors such as strain, mechanical perturbations[9], temperature and humidity fluctuations[10]. These environmentally induced variations can cause repetition frequency drift in the comb, leading to instability and inaccuracy in measurements[11][12]. The frequency instability, particularly the resulting jitter in the laser, can significantly degrade measurement resolution. Furthermore, absolute distance measurement experiments often need to be performed outdoors, where environmental control is limited[13]. Therefore, a highly integrated, robust and easy-tunable system which can stabilize the laser frequency is very advantageous and necessary in these applications.
Over the past few decades, various methods have been developed to stabilize pulse laser frequency. A common approach is the use of phase-locked loop (PLL) [14][15]. They can effectively lock the frequency of the laser to a specific value. However, cutting-edge experiments are often dynamic and complex, with changing requirements between measurements due to even minor differences in initial input parameters[16]. Traditional PLLs rely on analog electronics to track and correct phase shifts, but they are sensitive to environmental factors, such as cavity temperature fluctuations, making long-term stability difficult to maintain[14]. Additionally, analog components offer limited flexibility for adjustments. To address these limitations, digital PLLs (DPLLs) have been introduced, offering more flexibility and better resilience to environmental disturbances. However, they are still constrained by frequency dividers and multipliers, which hinder high-resolution frequency adjustments.
Currently, digital systems utilizing Micro-Controller Units (MCUs) or FPGAs are preferred for frequency-locked laser systems. These systems leverage their computational power to implement modern frequency measurement and feedback algorithms, achieving precise frequency locking. Digital systems measure frequency and provide feedback through ADCs and DACs, resulting in quantifiable error signals and adjustable feedback parameters that enhance flexibility and accuracy. While MCUs have been successfully employed in laser frequency stabilization using various locking algorithms[17][18], their feedback bandwidth is typically limited to the kHz range[19], restricting their performance in high-bandwidth applications.
To achieve high feedback bandwidth, FPGAs are an ideal platform for high-speed frequency locking circuits, thanks to their parallel processing capabilities and low latency[20]. Locking systems based on FPGAs can reduce system size while also offer high resolution and accuracy in data input and control output, handle high-concurrency data from high-speed ADC, provide real-time feedback processing, and support high-speed data transmission. Furthermore, modern FPGAs are programmable logic devices rich in computational resources, enabling the hardware implementation of complex algorithms required for laser frequency stabilization and easily configured to meet the requirements of different lasers.
The typical FPGA-based platforms like the Red Pitaya STEMlab have been increasingly adopted for real-world applications[21][22], including lock-in amplification, laser frequency stabilization, and intensity stabilization[23][24]. These platforms enable fully digital processing, which minimizes system noise while offering satisfactory timing resolution and high bandwidth[23][24]. However, lasers operate at various repetition frequencies, and fully digital methods face limitations due to lower ADC sampling rates. To prevent spectral aliasing, signals can only be captured below half the ADC’s sampling rate, which reduces adaptability to higher laser repetition frequencies. Simply increasing the ADC sampling rate is impractical, while it would compromise sampling accuracy. To address this issue, common approaches include frequency divider or mixing to lower the input signal frequency[8][10]. However, employing a frequency divider reduces the repetition frequency of the laser while also decreasing the error frequency that needs to be extracted, which leads to a loss in measurement accuracy. On the other hand, the mixing method necessitates the use of signal synthesizers and mixers, which may introduce additional noise.
@@
Figure 1: Schematic of locking circuits for laser frequency stabilization. PD: photodiode, BPF: band-pass filter, LPF: low-pass filter, CFM: configurable frequency multiplier, CFG: configurable frequency generator, PZT: piezoelectric transducer, HV: high voltage, Amp: voltage amplifier, ADC: analog-to-digital converter, DAC: digital-to-analog converter, DMTD: dual-mixer time-difference, PID: proportional integral derivative.
This paper presents an FPGA-based compact laser frequency locking method designed to adapt to various laser repetition frequencies while maintaining high precision and portability. The method utilizes error amplification and the dual-mixer time-difference (DMTD) method. The error amplification module (EAM) multiplies the frequency to amplify the error and subsequently down-converts it, reducing back-end hardware requirements. This design allows the system to adapt to various laser repetition frequencies while significantly enhancing measurement sensitivity with relatively low sampling rates. The method also integrates a digital frequency locking module (DFLM), include two high-speed, high-precision ADCs, a proportional-integral-derivative (PID) controller, and a high-precision DAC for feedback. All data processing algorithms are implemented on the FPGA, ensuring efficient real-time performance. System parameters are adjustable, allowing flexible measurement and frequency locking across a wide range of laser frequencies.
@@ II-A Overall Design
The schematic of the locking circuits for laser frequency stabilization is illustrated in Fig.1. This method measures the frequency difference between the laser repetition frequency and a high-stability reference source, an atomic clock in this case, enabling precise locking of the laser frequency. The pulse of the laser is first converted to electrical signal using a photodiode (PD). This signal is then processed by the EAM, which amplifies and down-converts the frequency error. The configurable frequency multiplier (CFM) performs frequency multiplication on the laser signal, producing high-frequency signals for mixing. The configurable frequency generator (CFG) uses the atomic clock to produce a stable high frequency signal also for mixing. After mixing, a low-pass filter (LPF) is used to obtain the amplified error signal. Through this process, the frequency error is amplified while the frequency of the signal to be sampled is lowered, thus reducing the requirements of the back-end circuitry.
@@
The error signal, alongside the atomic clock signal, is sampled by a high-speed ADC for processing by the FPGA. The FPGA utilizes the DMTD algorithm to calculate the phase difference between the sampled signals. This phase difference is converted into a frequency difference, passed through a low-pass filter, and then fed to the PID module, which controls a high-speed, high-precision DAC for feedback. Depending on the laser’s repetition frequency and the feedback devices employed, CFM, CFG and digital units within the FPGA can be configured to accommodate them. This capability facilitates the configuration of the EAM and the digital units within the FPGA, ensuring the method’s adaptability and robustness.
II-B Design of the Error Amplification Module (EAM)
@@
Advancements in technology have significantly improved the sampling rates and resolution of ADCs, making fully digital laser repetition frequency locking systems based on high-speed ADCs feasible[23]. In applications such as optical frequency combs, laser repetition frequencies can reach tens to hundreds of MHz, or even GHz[25]. However, typical fully digital laser locking systems operate with ADC sampling rates around 100 MHz, which, according to the Nyquist theorem, limits their ability to lock higher repetition frequency lasers. While utilizing higher sampling rates and resolutions could improve performance, the high costs make this impractical for many applications. To accommodate lasers with higher repetition frequencies, common techniques involve down-converting using methods such as frequency divider and mixing. Frequency divider lowers the frequency of the error signal during down-conversion, which can lead to precision loss, Mixing, on the other hand, requires signal synthesizers and mixers, which may introduce additional noise.
@@
To improve precision while reducing system costs, an EAM is designed to preprocesses the laser signal by incorporating principles of frequency multiplication and beat frequency methods. The PLL-based CFM multiplies the laser repetition frequency, while a PLL-based CFG, equipped with jitter-cleaning functionality, generates stable reference signals from the atomic clock for mixing and provides clock signals to the DFLM. Both the CFG and CFM feature zero-delay functionality, reducing system latency and increasing bandwidth.
While some approaches extract higher harmonics from the laser signal for mixing to obtain higher error frequencies, this requires BFP with narrow pass band, leading to poor out-of-band rejection, increased costs, and potential spectral aliasing, which may compromise measurement precision. Additionally, as the order of laser harmonics increases, the signal amplitude diminishes, and using high-gain operational amplifiers to boost the signal can introduce additional noise, making high-precision measurements more challenging.
By amplifying and down-converting the laser repetition frequency error using the EAM, measurement precision is enhanced while system costs are reduced, making the method more practical.
@@
Figure 2: Block diagram of measurement methods. (a) DQD method (b) DMTD method. REF: reference source.
II-C Design of DMTD and Frequency Extraction Module
In traditional laser repetition frequency locking systems, the digital quadrature demodulation (DQD) method, shown in Fig.2(a), is widely used for frequency measurement due to its simple structure and high resolution[20]. However, it is susceptible to inaccuracies caused by the NCO frequency generator. In contrast, DMTD method [26] can counteract noise from the common reference source, offering higher accuracy and flexible sampling period[27].
The analog DMTD method measures the phase difference of beat signals using zero-crossing detection and time interval counters[28][29]. This approach is prone to noise interference at the zero-crossing points, leading to measurement errors. Additionally, the presence of numerous analog components introduces extra noise into the system. Traditional DMTD methods control the measurement period by adjusting the beat frequency[30], which can reduce measurement resolution when the period is shortened.
This paper employs an ADC-based digital DMTD method Fig.2(b) to measure the frequency difference between two signals. This method avoids the zero-crossing noise issue associated with analog methods[31] and utilizes digital components such as mixers and filters. Digital components offer high consistency. In the symmetric design of the DMTD method, this consistency maximizes the cancellation of common-mode noise, significantly reducing noise and enhancing measurement accuracy. What’s more, this method does not require the two signals to have identical nominal frequencies, allowing for the measurement of signals with significant frequency deviations in the MHz range, which enhances the system’s measurement tolerance.
@@ II-C Design of DMTD and Frequency Extraction Module
${{U}_{r}}\left(t\right)={{A}_{r}}sin[2\pi{{f}_{r}}t+{{\varphi}_{r}}+\Delta{{\varphi}_{r}}(t)]$ (5)
@@
${{U}_{s2}}\left(t\right)={{A}_{s2}}cos\left(2\pi{{f}_{s2}}t+{{\varphi}_{s2}}+\Delta{{\varphi}_{s2}}(t)\right)$ (7)
@@
${{Q}_{1}}\left(t\right)\approx{{A}_{Q1}}\cos\left[2\pi{{f}_{h1}}t-{{\varphi}_{s1}}+{{\varphi}_{r}}-\Delta{{\varphi}_{s1}}(t)+\Delta{{\varphi}_{r}}(t)\right]$
@@
By subtracting ${{\theta}_{2}}(t)$ from ${{\theta}_{1}}(t)$, the phase difference between the amplified error frequency signal and the reference signal is calculated as:
@@
$=2\pi\left({{f}_{r}}-{{f}_{e}}\right)t+{{\varphi}_{r}}-{{\varphi}_{e}}+\Delta{{\varphi}_{r}}\left(t\right)-\Delta{{\varphi}_{e}}\left(t\right)$
@@
$=m\Delta f+\frac{\Delta{{\varphi}_{r}}{{\left(t\right)}^{\prime}}-\Delta{{\varphi}_{e}}{{\left(t\right)}^{\prime}}}{2\pi}$
This frequency difference includes high-frequency phase noise $\frac{\Delta{{\varphi}_{r}}{{\left(t\right)}^{\prime}}-\Delta{{\varphi}_{e}}{{\left(t\right)}^{\prime}}}{2\pi}$, which introduces measurement errors. However, the frequency difference $m\Delta f$ remains relatively stable over short periods, while the phase noise decreases when averaged. Thus, by averaging over short periods, the phase noise decreases, and the frequency difference stabilizes:
${\overline{{f}_{r,e}}}=m\Delta f+\overline{\frac{\Delta{{\varphi}_{r}}{{\left(t\right)}^{\prime}}-\Delta{{\varphi}_{e}}{{\left(t\right)}^{\prime}}}{2\pi}}$ (16)
With sufficient averaging, the phase noise diminishes, and the frequency difference approaches $m$, providing a stable input for digital feedback algorithms that lock the laser repetition frequency.
@@
Due to inevitable differences between the NCO signal, the measured signal, and the reference signal, the phase between them slowly but steadily drifts, causing phase wraps. During frequency calculation, phase unwrapping is required to obtain accurate data. Differentiating the phase data can convert it into instantaneous frequency. Then averaging the frequency difference can reduce the impact of phase noise.
Additionally, different feedback devices operate at varying response speeds, and it may be necessary to adjust the measurement period to meet specific requirements. To address these challenges and minimize phase noise, a counter-based phase unwrapping method is employed for frequency calculation and decimation. This method requires only a counter and minimal RAM, allowing for frequency averaging over a wide range. Fig.3 illustrates the phase unwrapping process, where the counter increments or decrements depending on the direction of the phase wrap. By adding the counter’s phase value to the current phase and differentiating over different time intervals, frequency measurements can be obtained.
In a digital system, signals consist of discrete sampled points. Based on the derived approach above, a series of phase points is obtained within the FPGA, denoted as ${\theta(1),\theta(2),\ldots,\theta(n-1),\theta(n)}$. The frequency at any given point is expressed as:
$f(n)=\frac{\theta(n)-\theta(n-1)}{2\pi}.$ (17)
@@
$\overline{f}=\frac{1}{2\pi m}\sum_{n=2}^{m}\left(\theta(n)-\theta(n-1)\right),$ (18)
where $m$ is the length of the measurement period. Simplifying this expression yields:
$\overline{f}=\frac{\theta(m)-\theta(1)}{2\pi m}.$ (19)
Thus, only the first and last phase values need to be stored in the FPGA to calculate the frequency over the cycle of $m$. Using a counter, the unwrapped phase values $\theta^{\prime}(1)$ and $\theta^{\prime}(m)$ can be determined as follows:
$\theta^{\prime}(1)=\theta(1),$ (20)
$\theta^{\prime}(m)=\theta(m)+2\pi\text{CNT}(m).$ (21)
@@
$\overline{f}=\frac{\theta^{\prime}(m)-\theta^{\prime}(1)}{2\pi m}.$ (22)

### v2

# A Digital and Compact High-Precision Locking System for Pulse Laser Repetition Frequency

@@ Abstract
This paper proposes a novel approach that employs error amplification and ADC-based dual-mixer time-difference (ADC-based-DMTD) technique for high-precision locking of laser repetition frequency with compact size. This electronic system consists of two main components: a digitized error amplification module (EAM) and an FPGA-based digital frequency locking module (DFLM). The EAM mainly integrates a configurable frequency generator (CFG), a configurable frequency multiplier (CFM) and a mixer to process the laser pulses and a high-stability reference source (e.g., an atomic clock), enabling high-precision locking of pulse lasers operating at different repetition frequencies. By employing frequency multiplication and mixing, the EAM amplifies the laser’s frequency error and performs frequency down-conversion, enhancing measurement sensitivity and reducing the hardware requirements of the back-end. The DFLM receives the EAM outputs by using an ADC-based-DMTD method to precisely measure frequency errors, then the digital proportional-integral-derivative (PID) controller provides feedback to achieve accurate frequency locking. Initial testing with a voltage-controlled oscillator (VCO) demonstrated excellent locking performance, achieving an Allan deviation of $9.58\times 10^{-14}$ at 10 seconds and a standard deviation (STD) of 7.7 µHz root mean square (RMS) after locking, marking a five-order-of-magnitude stability enhancement. In laboratory experiments with a custom-built femtosecond fiber laser, the system achieved robust locking of the repetition frequency, with a stability improvement from $1.51\times 10^{-7}$ to $1.12\times 10^{-12}$ at a 10-second gate time and an STD of 0.43 mHz RMS after locking.
@@
Field programmable gate array (FPGA), laser frequency locking, frequency error amplifier, dual-mixer time-difference (DMTD).
@@ I Introduction
Pulse lasers characterized by narrow-lines and high repetition frequency stability are essential components in optical frequency combs. They have widespread applications in both industrial settings and scientific research, including high-precision measurement[1], [2], optical communication[3], [4], high-resolution optical spectroscopy[5], [6], etc. Optical frequency combs are advanced tools that generate millions of mutually coherent and evenly spaced modes, enabling precise phase-coherent connections across the optical, microwave, and terahertz frequency ranges. These frequency combs are essential for a variety of applications and fundamental scientific research[7], [8]. However, optic combs have been limited by various factors such as strain, mechanical perturbations[9], temperature and humidity fluctuations[10]. These environmentally induced variations can cause repetition frequency drift in the comb, leading to instability and inaccuracy in measurements[11], [12]. The frequency instability, particularly the resulting jitter in the laser, can significantly degrade measurement resolution. Furthermore, many experiments often need to be performed outdoors, where environmental control is limited, and system size is strictly constrained[13], [14], [15]. Therefore, a highly integrated, robust and easy-tunable system which can stabilize the laser repetition frequency with high precession is very advantageous and necessary in these applications.
Methods such as analog phase-locked loop (PLL) and charge pump PLL (CP-PLL) have been employed for laser repetition frequency stabilization[16], [17] in past few decades. However, these methods are constrained by environmental sensitivity, feedback accuracy, and limited tunability. Since laser experiments are often dynamic and complex, achieving precise and flexible frequency locking with configurable parameters necessitates the use of advanced digital systems. Digital systems based on microcontroller units (MCUs) and FPGAs have emerged as preferred approaches for laser frequency stabilization[18], [19]. While MCUs typically offer bandwidths in the kHz range[20], FPGAs provide higher bandwidth and are better suited for applications requiring greater precision and flexibility[21].
The typical FPGA-based platforms like the Red Pitaya STEMlab have been increasingly adopted for real-world applications[22], [23], [24]. Some systems use fully digital FPGA architectures for frequency measurement and feedback [25], [26], [27], which eliminates interference from analog components while offering excellent timing resolution and high bandwidth. But lasers typically operate at repetition frequencies in the tens to hundreds of MHz, or even GHz [28], making direct signal capture with ultra-high-speed ADCs prohibitively expensive. To address this issue, common approaches include frequency divider or higher-order harmonics mixing to lower the input signal frequency[8], [10]. Frequency dividers reduce the repetition frequency error, leading to a loss in measurement accuracy. The higher-order harmonics mixing method offering a solution for improving measurement accuracy for it contains greater frequency error as the higher harmonics of the laser increase. However, the accompanying reduction in energy results in a lower signal-to-noise ratio, making them highly susceptible to environmental influences.
@@
Figure 1: Schematic of locking circuits for laser frequency stabilization. PD: photodiode, BPF: band-pass filter, LPF: low-pass filter, CFM: configurable frequency multiplier, CFG: configurable frequency generator, PZT: piezoelectric transducer, HV: high voltage, Amp: voltage amplifier, ADC: analog-to-digital converter, DMTD: dual-mixer time-difference, DAC: digital-to-analog converter, PID: proportional integral derivative.
In this paper, we present an FPGA-based, high-precision system for laser repetition frequency locking. The system employs digital error amplification and ADC-based dual-mixer time-difference (ADC-based-DMTD) technique to achieve accurate measurement. The digital error amplification module (EAM) uses digital PLL chip to multiply the repetition frequency, which multiplies the frequency errors while down-converting the signal to simplify the requirements of back-end hardware. The approach enhances sensitivity and provides flexibility for adapting to different repetition frequencies. The digital frequency locking module (DFLM) captures the high-stability reference source and error signals via high-speed, high-precision ADCs. By referencing the DMTD method [29], [30], DFLM utilizes an ADC-based-DMTD method to perform high-precision phase measurements. Followed by frequency extraction and filtering, this processed signal is sent to a PID controller, which adjusts the DAC to achieve real-time feedback, ensuring precise laser frequency locking and long-term stability. Both the EAM and DFLM are closely linked with the FPGA, allowing for easy system parameter adjustments to adapt to different laser repetition frequencies. By utilizing these methods, the system improves measurement accuracy while maintaining a highly integrated structure, and can provide robust laser stabilization across a wide range of repetition frequencies, suitable for diverse applications.
The paper is organized as follows: Section II gives an overview of the laser repetition frequency locking system, detailing the system design and theory analysis. Section III describes the implementation of the electronic prototype and presents performance evaluations through noise floor testing and VCO stability experiments. A laser stability experiment is conducted to analyze long-term frequency fluctuations and Allan deviation before and after locking. A summary of this research and the discussion of future improvements are given in Section IV.
@@ II-A Overall Design
The schematic of the locking circuits for laser frequency stabilization is illustrated in Fig.1. This system measures the frequency difference between the laser repetition frequency and a high-stability reference source, an atomic clock in this case, enabling precise locking of the laser frequency. The pulse of the laser is first converted to electrical signal using a photodiode (PD). This signal is then processed by the EAM, which amplifies and down-converts the frequency error. The configurable frequency multiplier (CFM) performs frequency multiplication on the laser signal, producing high-frequency signals for mixing. The configurable frequency generator (CFG) uses the atomic clock to produce a stable high frequency signal also for mixing. After mixing, a low-pass filter (LPF) is used to obtain the amplified error signal. Through this process, the frequency error is amplified while the frequency of the signal to be sampled is lowered, thus reducing the requirements of the back-end circuitry.
@@
The error signal and atomic clock signal are sampled by high-speed ADCs for FPGA processing. An ADC-based-DMTD algorithm is adopted to calculate the phase difference between the sampled signals. Then this phase difference is converted into a frequency difference, passed through a low-pass filter, and fed to the PID module, which controls a high-speed, high-precision DAC for feedback. Depending on the laser’s repetition frequency and the feedback devices employed, CFM, CFG and digital units within the FPGA can be configured to accommodate them. This capability facilitates the configuration of the EAM and the digital units within the FPGA, ensuring the system’s adaptability and robustness.
II-B Design of the digitized Error Amplification Module (EAM)
@@
Advancements in technology have significantly improved the sampling rates and resolution of ADCs, making fully digital laser repetition frequency locking systems based on high-speed ADCs feasible[24]. In applications such as optical frequency combs, laser repetition frequencies can reach tens to hundreds of MHz, or even GHz[28]. While ultra-high-speed ADCs can directly capture these signals, this approach is costly and introduces considerable design complexity. To accommodate lasers with higher repetition frequencies, common techniques involve down-converting using methods such as frequency divider and high-order harmonic mixing[8], [10]. Frequency divider lowers the frequency of the error signal during down-conversion, which can lead to precision loss. High-order harmonic mixing improves accuracy by extracting higher harmonics of the laser signal, which contains higher frequency error signal. But the narrow pass band BFP used to extract higher harmonics often suffer from poor out-of-band rejection, increased costs, and potential spectral aliasing, all of which can compromise measurement accuracy. Additionally, as harmonic order increases, signal amplitude decreases, necessitating high-gain amplification. This amplification, while boosting the signal, introduces noise, making high-precision measurements more challenging.
@@
To improve precision while reducing hardware costs, a digitized EAM is designed to preprocesses the laser signal by incorporating principles of frequency multiplication and beat frequency methods. The PLL-based CFM multiplies the laser repetition frequency, while a PLL-based CFG, equipped with jitter-cleaning functionality, generates stable reference signals from the atomic clock for mixing and provides clock signals to the DFLM. The digital method for frequency multiplies ensures the energy of high-frequency signals while providing strong resistance to interference. By amplifying and down-converting the laser repetition frequency error using the digitized EAM, measurement precision is enhanced while hardware costs are reduced.
@@
Figure 2: Block diagram of measurement methods. (a) digital quadrature demodulation (DQD) method (b) ADC-based-DMTD method. REF: reference source, NCO: numerically controlled oscillator.
II-C Design of ADC-based-DMTD and Frequency Extraction Module
In traditional laser repetition frequency locking systems, the digital quadrature demodulation (DQD) method [31], [8], shown in Fig.2(a), is widely used for frequency measurement due to its simple structure and easy-tunable ability. However, it is susceptible to inaccuracies caused by the instability and phase noise from the numerically controlled oscillator (NCO). In contrast, DMTD method[29], [30] can counteract noise from the NCO, offering higher accuracy while maintaining flexible sampling period[32].
The traditional analog DMTD method measures the phase difference of beat signals through zero-crossing detection and time interval counters [33], [34]. However, the reliance on analog components introduces additional noise into the system [33]. While the digital DMTD (D-DMTD) [35] approach eliminates analog noise by using fully digital techniques, both analog and digital methods are susceptible to noise interference at zero-crossing points, leading to measurement errors. Additionally, these methods adjust the measurement period by altering the beat frequency [36], which can lower resolution when the measurement period is shortened.
This paper employs an ADC-based-DMTD method Fig.2(b) to measure the frequency difference between two signals. This method avoids the zero-crossing noise issue associated with analog methods and utilizes digital components, such as mixers and filters, which offer greater consistency[37], [38]. In the symmetric design of the ADC-based-DMTD method, this consistency maximizes the cancellation of common-mode noise, significantly reducing noise and enhancing measurement accuracy. Additionally, the method does not require the two signals to have identical nominal frequencies, enabling it to measure signals with substantial frequency deviations in the MHz range. This feature broadens the system’s measurement tolerance.
@@ II-C Design of ADC-based-DMTD and Frequency Extraction Module
${{U}_{r}}\left(t\right)={{A}_{r}}sin[2\pi{{f}_{r}}t+{{\varphi}_{r}}+\Delta{{\varphi}_{r}}(t)].$ (5)
@@
${{U}_{s2}}\left(t\right)={{A}_{s2}}cos\left(2\pi{{f}_{s2}}t+{{\varphi}_{s2}}+\Delta{{\varphi}_{s2}}(t)\right).$ (7)
@@
[…]
@@
By subtracting ${{\theta}_{2}}(t)$ from ${{\theta}_{1}}(t)$, the common-mode noise $\Delta{{\varphi}_{s1}}\left(t\right)$ from the NCO was canceled. The phase difference between the amplified error frequency signal and the reference signal is calculated as:
@@
$=2\pi\left({{f}_{r}}-{{f}_{e}}\right)t+{{\varphi}_{r}}-{{\varphi}_{e}}+\Delta{{\varphi}_{r}}\left(t\right)-\Delta{{\varphi}_{e}}\left(t\right).$
@@
$=m\Delta f+\frac{\Delta{{\varphi}_{r}}{{\left(t\right)}^{\prime}}-\Delta{{\varphi}_{e}}{{\left(t\right)}^{\prime}}}{2\pi}.$
In traditional DQD method, the phase noise from the NCO is still existed, but it is counteracted in DMTD method. This frequency difference ${f}_{r,e}$ includes high-frequency phase noise $\frac{\Delta{{\varphi}_{r}}{{\left(t\right)}^{\prime}}-\Delta{{\varphi}_{e}}{{\left(t\right)}^{\prime}}}{2\pi}$, which introduces measurement errors. However, the frequency difference $m\Delta f$ remains relatively stable over short periods, while the phase noise decreases when averaged. Thus, by averaging over short periods, the phase noise decreases, and the frequency difference stabilizes:
${\overline{{f}_{r,e}}}=m\Delta f+\overline{\frac{\Delta{{\varphi}_{r}}{{\left(t\right)}^{\prime}}-\Delta{{\varphi}_{e}}{{\left(t\right)}^{\prime}}}{2\pi}}.$ (16)
With averaging, the phase noise diminishes, and the frequency difference approaches $m$, providing a stable input for digital feedback algorithms that lock the laser repetition frequency.
Due to inevitable differences between the NCO signal, the measured signal, and the reference signal, the phase between them slowly but steadily drifts, causing phase wraps[39]. During frequency calculation, phase unwrapping is required to obtain accurate data. Differentiating the phase data can convert it into instantaneous frequency. Then averaging the frequency difference can reduce the impact of phase noise.
@@
Additionally, different feedback devices operate at varying response speeds, and it may be necessary to adjust the measurement period to meet specific requirements. To address these challenges and minimize phase noise, a counter-based phase unwrapping method is employed for frequency calculation and decimation. This method requires only a counter (CNT) and minimal RAM, allowing for frequency averaging over a wide range. Fig.3 illustrates the phase unwrapping process, where the CNT increments or decrements depending on the direction of the phase wrap. By adding the CNT’s phase value to the current phase and differentiating over different time intervals, frequency measurements can be obtained.
In a digital system, signals consist of discrete sampled points. As the sample rate of the ADC is $F_{s}$, its sample period is $T_{s}$. Based on the derived approach above, a series of phase points is obtained within the FPGA, denoted as ${\theta(1),\theta(2),\ldots,\theta(n-1),\theta(n)}$. The frequency at any given point is expressed as:
$f(n)=\frac{\theta(n)-\theta(n-1)}{2\pi T_{s}}.$ (17)
@@
$\overline{f}=\frac{1}{2\pi(m-1)T_{s}}\sum_{n=2}^{m}\left(\theta(n)-\theta(n-1)\right)$ (18)
where $m-1$ is the length of the measurement period. Simplifying this expression yields:
$\overline{f}=\frac{\theta(m)-\theta(1)}{2\pi(m-1)T_{s}}.$ (19)
Thus, only the first and last phase values need to be stored in the FPGA to calculate the frequency over the cycle of $m$. Using a counter, the unwrapped phase values $\theta_{u}(1)$ and $\theta_{u}(m)$ can be determined as follows:
$\theta_{u}(1)=\theta(1)$ (20)
$\theta_{u}(m)=\theta(m)+2\pi\text{CNT}(m).$ (21)
@@
$\overline{f}=\frac{\theta_{u}(m)-\theta_{u}(1)}{2\pi(m-1)T_{s}}=\frac{\theta(m)-\theta(1)+2\pi\text{CNT}(m)}{2\pi(m-1)T_{s}}.$ (22)